OpenStudy (rahulmr):

the solution to \[b*\sin(ax)=\sqrt{2}*\cos(ax)\] is \[\frac{ \pi }{ 4 }\]. What are the possible values of a,b. Help anyone. I don't know how should i solve this. Thanks.

zepdrix (zepdrix):

Wow you have some weird tricky problems lol

OpenStudy (rahulmr):

\[b*\tan(ax)=\sqrt{2}\] i think it will simplify to this.

OpenStudy (rahulmr):

yes. it's my homework.

zepdrix (zepdrix):

\[\large\rm \tan(ax)=\frac{\sqrt2}{b}\]I'm still trying to figure this one out. But here is what I'm thinking so far... By plugging in key values for a, you can come up with corresponding b values that work. For example: a=1, well tangent of pi/4 is 1, ya? So b=sqrt2.

zepdrix (zepdrix):

a cannot equal 0, that would cause a problem for b. a cannot equal 2 as tangent of pi/2 is not defined. But then we get all these weird numbers in between... When a=4/3, tangent of pi/3 gives us sqrt3, so b= ummm.. like sqrt2/sqrt3, ya?

OpenStudy (rahulmr):

yes. you are right a=1 and b=sqrt{2}

OpenStudy (rahulmr):

if i sub these values in a calculator, it gives me somewhat pi/4

OpenStudy (rahulmr):

You are a genius. Thanks for the help again.

zepdrix (zepdrix):

I'm not sure how we get the entire collection of a and b's, hmm :\

OpenStudy (rahulmr):

don't worry, i just need one possible value

zepdrix (zepdrix):

Just one? Oh interesting :D

OpenStudy (rahulmr):

yeah !!